Problem 1: A Newtonian fluid is confined between the plate and a fixed surface. If the velocity profile is defined by u =(8y- answer is 4.05(10^-5) N
Answer #1

Here le= 0.482 x 100% N-s. ma - 0:482 xlo? N -s mma los mma = 0.482 x 10 9 N -s Now that by osy 1. u= sy. 0.3y² du =8-0.by at